# quality of service 5G fifth generation

### v1

@@
Physical Resource Blocks
@@
Salar ${}^{\dagger}$    Mojdeh Karbalaee ${}^{\dagger}$     Vahid ${}^{\dagger}$       Seyed Pooya ${}^{\dagger}$
@@ Abstract
The Open Radio Access Network (O-RAN) technology has emerged as a promising solution for network operators, providing them with an open and favorable environment. Ensuring effective coordination of x-applications (xAPPs) is crucial to enhance flexibility and optimize network performance within the O-RAN. In this paper, we introduce an innovative approach to the resource allocation problem, aiming to coordinate multiple independent xAPPs for network slicing and resource allocation in O-RAN. Our proposed method focuses on maximizing the weighted throughput among user equipments (UE), as well as allocating physical resource blocks (PRBs). We prioritize two service types, namely enhanced Mobile Broadband and Ultra Reliable Low Latency Communication. To achieve this, we have designed two xAPPs: a power control xAPP for each UE and a PRB allocation xAPP. The proposed method consists of a two-part training phase, where the first part uses supervised learning with a Variational Autoencoder trained to regress the power transmission as well as the user association and PRB allocation decisions, and the second part uses unsupervised learning with a contrastive loss approach to improve the generalization and robustness of the model. We evaluate the performance of our proposed method by comparing its results to those obtained from an exhaustive search algorithm, deep Q-network algorithm, and by reporting performance metrics for the regression task. We also evaluate the proposed model’s performance in different scenarios among the service types. The results show that the proposed method is a more efficient and effective solution for network slicing problems compared to state-of-the-art methods.
@@ I INTRODUCTION
Wireless communication has evolved from a mere convenience to an absolute necessity, driven by an insatiable demand for faster and more dependable connectivity between various quality of service (QoS) requirements [1]. The emergence of fifth generation (5G) technology represents a significant leap, offering the potential to support a wide range of applications, including augmented reality and autonomous vehicles [2]. However, the pursuit of wireless excellence does not conclude with 5G; it merely marks a stepping stone towards the next frontier— sixth generation (6G), the sixth generation of wireless technology. Poised on the horizon, 6G holds the promise to once again redefine the boundaries of connectivity.
@@
Emerging concepts within 6G, such as extreme Multiple-Input Multiple-Output and deep Receiver, promise to propel us further into uncharted technological realms, offering unprecedented data rates and levels of connectivity [3]. These advancements have the potential to not only revolutionize our modes of communication but also reshape the very fabric of our interconnected world. Furthermore, the application of various machine learning (ML) techniques to automate resource allocation within the medium access control (MAC) layer and scheduler represents a burgeoning area of research within the 6G landscape [4]. These techniques aim to inject an unparalleled level of intelligence into network management and optimization, laying the foundation for networks capable of dynamically adapting to user demands and environmental conditions.
In this constantly evolving landscape, the roles of Open Radio Access Network (O-RAN) [5] and network slicing [6], which have reached remarkable milestones in 5G, continue to be pivotal. As we explore the possibilities offered by 6G and its transformative capabilities, it becomes increasingly evident that the evolution of wireless communication is an unceasing journey. Each generation builds upon the accomplishments of its forerunners, establishing a robust foundation for the next phase of wireless excellence.
@@
Resource allocation involves the meticulous distribution of finite network resources among numerous users, applications, or services [7]. In 5G, the primary objective is to support a wide spectrum of services, including Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communications (URLLC), and massive Machine Type Communications (mMTC) [8]. Each of these services necessitates distinct resource sets encompassing bandwidth, data rate, latency, power consumption, complexity, and reliability. As such, resource allocation within the 5G landscape must exhibit the flexibility to cater to these diverse demands while ensuring the efficient utilization of network resources.
@@
Network slicing is an innovative concept that tailors virtual networks (slices) within a unified physical infrastructure, addressing specific service types and use cases [9]. In the context of 5G, it efficiently allocates resources for eMBB and URLLC [10] services. This significance extends into the anticipated 6G era, where network slicing remains instrumental, facilitating resource allocation for diverse applications, including holographic communication and tactile internet [11, 12]. Its inherent scalability ensures adaptability to evolving workloads and user expectations, solidifying its role as a key technology in the relentless pursuit of wireless excellence, bridging the transition from 5G to 6G.
@@
This concept seamlessly aligns with the architecture of the O-RAN. O-RAN is based on open interfaces, open hardware, and open software and aims to provide a more flexible and cost-effective solution for wireless communication [7]. Within this context, the implementation of network slicing in O-RAN not only optimizes resource allocation but also fosters multi-vendor interoperability and innovation.
@@
At the core of O-RAN’s architecture are two fundamental components: Radio Units, responsible for physical layer functions, and Distributed Units, which manage baseband processing and control. These components collaborate seamlessly, with the Central Unit (CU) serving as a centralized coordinator for network resources, facilitating dynamic resource allocation [13, 5]. The CU further comprises two distinct elements: the control plane (CP) and the user plane (UP). Adding to O-RAN’s adaptability and innovation are xApps (cross-applications), capable of spanning multiple network domains to optimize resource allocation for various services. These network segments include Fronthaul, Backhaul, and Midhaul, facilitating data transmission between components. Overseeing resource management and orchestration tasks is the Management and Orchestration (MANO) layer, ensuring the network’s efficiency. Network Functions Virtualization (NFV) decouples software from hardware, enhancing scalability. Furthermore, the Radio Access Network Intelligent Controller (RIC) plays a pivotal role in intelligent resource control, enabling dynamic allocation to meet diverse service requirements [5]. These components collectively define O-RAN’s transformative architecture, promoting openness, adaptability, and efficient resource utilization in modern wireless networks. It is worth noting that O-RAN comprises two types of RIC: near-real-time (near-RT) RIC, which manages Radio Access Network (RAN) control and optimization, and non-real-time (non-RT) RIC. Within the near-RT RIC, xAPPs facilitate RAN control and optimization, while inside the non-RT RIC, rAPPs support it by providing guidance and employing ML models. Furthermore, both xAPPs and rAPPs can be implemented by third-party entities, allowing for a diverse ecosystem of applications within the O-RAN architecture [14].
@@
Moreover, xAPPs leverage O-RAN network slicing for efficient resource utilization, allowing seamless collaboration across services. However, this introduces network management complexity, as each xAPP must adapt to multiple slice QoS needs, posing a critical coordination challenge. Additionally, cellular operators face the difficulty of establishing consistent connectivity for devices with different QoS demands [15, 16]. Therefore, in this study, we examine the resource allocation problem using network slicing by considering the xAPPs. We propose a novel method for resource allocation, with the goal of effectively coordinating numerous independent xAPPs to address the network slicing and resource allocation challenges within O-RAN.
The application of ML algorithms in wireless networks has surged due to the escalating complexity of these networks and the challenges posed by crafting custom optimization models. Traditional approaches, such as convex/non-convex optimization, are grappling with mounting complexity stemming from shifts in network architecture, evolving performance requirements, and the proliferation of devices. Present methods for network resource allocation that rely on learning-based techniques, like deep learning (DL) and deep reinforcement learning (DRL), often neglect the capacity to learn resource distributions. Typically, these methods involve designing an agent for a specific task and training it from scratch, resulting in reduced exploration efficiency and extended convergence times. Moreover, they necessitate a substantial volume of training samples, leading to prolonged exploration phases. The limited adaptability of traditional ML methods to novel scenarios motivates us to explore an alternative approach capable of superior generalization, more efficient data utilization, and resource distribution learning. To address these objectives, we propose a novel methodology that employs a self-supervised learning technique, specifically contrastive learning [17], in conjunction with a variational autoencoder (VAE) model, known as a type of generative model [18]. These models excel in handling uncertainty by learning potential outcome distributions, thereby offering a distribution of predictions instead of a singular deterministic output. Moreover, their adeptness in handling incomplete or missing data allows for the imputation of missed values based on learned data patterns. VAEs also exhibit proficiency in semi-supervised learning setups, effectively harnessing both labeled and unlabeled data to enhance predictive accuracy. Their generative abilities facilitate the creation of new data instances resembling the training set, proving beneficial in scenarios demanding synthetic data generation or heightened resilience against noise and outliers.
@@
Furthermore, contrastive learning involves contrasting different examples to acquire informative data representations. On the other hand, a VAE comprises an encoder and decoder that jointly learn the underlying probability distribution of data. By integrating contrastive learning with VAE, we can mitigate the common challenge of requiring a large number of training samples in ML. This approach can enhance performance metrics, such as quicker convergence and improved results, while also reducing the time and effort needed for algorithm design and training.
@@ I-A Contribution
This paper introduces a system model within the O-RAN architecture that provides support for two primary services, namely eMBB and URLLC, through resource allocation, with a particular emphasis on network slicing.
Our paper’s primary contributions can be summarized as follows:
  • Our research primarily aims to tackle resource allocation challenges in the O-RAN architecture, with a specific focus on optimizing weighted throughput by allocating resources like transmission power and physical resource blocks. We account for the unique QoS requirements and service priorities of two key 5G services: eMBB and URLLC. Compared to existing joint resource allocation schemes, our xAPP-based architecture enhances network management flexibility.
  • Our solution to the resource allocation optimization problem employs a two-step algorithm. Firstly, we simplify the problem by framing it as an optimization task. Secondly, we employ a two-phase training approach. We use supervised learning with a VAE to train power transmission regression and resource allocation decisions. This is followed by unsupervised learning with a contrastive loss method to enhance the model’s generalization and robustness. In essence, our approach simultaneously handles power and radio resource allocation xAPPs. In addition, it is not limited to O-RAN but can be applied to future multi-vendor O-RAN architectures, leveraging team learning principles developed by our team.
@@ I-B Structure of Paper
The paper is structured as follows: Section II presents a review of related works. In Section III, we introduce the system model for the resource allocation problem in the O-RAN architecture and formally define the resource allocation problem as an optimization problem. Section IV provides a detailed description of the proposed resource allocation algorithm. In Section V, we present numerical results used to evaluate the algorithm’s performance. Finally, Section VI offers our conclusion.
@@ II Related Works
As mentioned in Section I, cellular operators face significant challenges in ensuring reliable connections for devices with diverse QoS requirements. This challenge becomes more pronounced in 5G networks due to their increased complexity and stringent QoS demands [16]. Recently, there has been a growing focus on addressing the network slicing problem in cellular networks, as evidenced by several studies [16, 19, 20, 21, 22].
Feng et al. [16] proposed a method that employs network slicing and scheduling to allocate resources based on QoS requirements. However, their simulation utilized only a limited number of User Equipment (UE) devices and service types, which may not accurately represent real-world 5G networks with diverse service requirements. In another study, a resource allocation approach based on the M/M/1 queuing system was presented for RAN slicing [19]. Nevertheless, practical scenarios often involve bursty traffic, rendering the M/M/1 assumption of a Poisson distribution of packet arrivals at each gNodeB, which serves the eMBB and URLLC services, overly optimistic. Consequently, the gNodeB cannot be considered an M/M/1 system, and its queue cannot be accurately modeled as a continuous-time Markov process. Additionally, in a study by Kazemifard et al. [20], an O-RAN system was investigated, assuming that the core network and RAN are co-located in the same data center. The authors conducted a theoretical analysis and employed a mathematical model to efficiently address minimum delay problems.
Various approaches have been proposed in the literature to address resource allocation challenges in wireless networks, including optimization techniques, game theory, and ML algorithms [23]. ML algorithms, in particular, have found application in power control and resource allocation within wireless networks. These models are trained on historical resource allocation decisions and then applied to make new allocation decisions.
@@
Similarly, Samarakoon et al. [30] introduced a federated learning algorithm to optimize transmit power and radio resource allocation in vehicular networks. Additionally, Elsayed et al. [31] proposed a reinforcement learning-based method for joint power and radio resource allocation in 5G networks.
@@ III System Model and Problem Formulation
In our study, we address network slicing and resource allocation problem within O-RAN architecture, focusing on two primary slices: the URLLC slice and the eMBB slice. The URLLC slices ensure ultra-low latency and high reliability for critical applications such as autonomous driving and remote surgery, while eMBB slices focus on high data rates for mobile broadband. These slices have also various QoS requirements for network resource allocation [10].
By slicing the RAN into dedicated components for URLLC and eMBB, we can tailor network optimization to distinct traffic types and services, enhancing the efficient and flexible utilization of network resources. Key resources in O-RAN allocation encompass transmission power and Physical Resource Blockss. Traffic demands, often quantified by data rate and bandwidth, reflect the volume of data requiring transmission through the network. QoS requirements delineate the desired network performance based on parameters such as data rate. Our system model for resource allocation with network slicing is illustrated in Figure 1.
@@
In our study, we investigate an O-RAN-based deployment, wherein each base station, known as RU in O-RAN, accommodates two distinct slices: eMBB and URLLC. These slices cater to UEs with comparable QoS requirements. Our analysis encompasses two critical network functions: power control and radio resource allocation, which are jointly considered. We denote the set of all RUs as $\mathcal{B}$ and the set of UEs as $\mathcal{U}$. Each base station $b$ serves two sets of users: $\mathcal{U}_{e,b}$ for eMBB and $\mathcal{U}_{u,b}$ for URLLC, leading to a total of $U$ users, where $U=U_{e,b}+U_{u,b}$. Each user is exclusively associated with a single RU. The temporal and spectral layout is partitioned into $\mathcal{M}$ segments referred to as PRBs. These PRBs constitute the smallest unit of resources for allocation [32]. Each PRB spans a time frame, defined as a time slot in the temporal domain, and covers a frequency range of $f$ Hz in the spectral domain. We also denote $S_{e}$ for eMBB slices and $S_{u}$ for URLLC slices. The total number of slices ($S$) will be: $S=S_{e}+S_{u}$.
The user association is indicated by a binary variable $\alpha_{u,b,s}$, which takes a value of 1 if user $u$ is connected to RU $b$ in slice $s$, and 0 if not. Since each UE should be connected to only one RU, the following constraint must be met:
@@
We define the binary variable $\beta_{u,b,m,s}$ as the indicator for PRB allocation, taking the value of 1 when user $u$ in slice $s$ is connected to PRB $m$ of RU $b$, and 0 otherwise.
@@
The rate achieved for a user $u$ by allocating PRB $m$ of RU $b$ is denoted by $R_{u,b,m,s}$, where $\eta_{u,b,m,s}$ represents the Signal to Interference & Noise Ratio (SINR). The calculation for $R_{u,b,m,s}$ is determined by:
@@
The total power transmitted by RU $b$ and the data rate for UE in the fronthaul link are represented as $P_{b}$ and $C_{b}$, respectively. The existence of quantization noise ($\sigma_{1}^{2}$) can potentially impact the signal’s fidelity during transmission over the fronthaul, particularly when the original signal displays significant variations or spans a wide range of values. This noise may lead to inaccuracies in the transmitted signal, thereby influencing the QoS perceived by the UE. These values are determined as follows:
@@
where, we denote the transmission power of PRB $m$ in RU $b$ in slice s as $p_{u,b,m,s}$ for UE $u$, and the channel gain between RU $b$ and UE $u$ in slice $s$ as $h_{u,b,m,s}$. Furthermore, we define the channel gain between the resources and UE as $h_{ij}$. To represent this relationship in matrix form, we construct a matrix that accounts for $\mathcal{B}$ resources and $\mathcal{U}$ users as follows:
$H:=\begin{bmatrix}|H_{11}|^{2}&\cdots&|H_{1U}|^{2}\\ \vdots&\vdots&\vdots\\ |H_{B1}|^{2}&\cdots&|H_{BU}|^{2}\end{bmatrix}_{B\times U}$ (13)
The overall propagation delay ($D^{pro}$) comprises three distinct components: the propagation delay in the fronthaul link ($D^{fr,p}$), the midhaul link ($D^{mid,p}$), and the backhaul link ($D^{b,p}$). In each of these links, the propagation delay is determined by the time it takes for a signal to traverse the distance, calculated as $D=\frac{L}{C}$, where $L$ signifies the length of the fiber link, and $C$ represents the propagation speed of the medium.
Similarly, the total transmission delay ($D^{tr}$) is the aggregate of transmission delays in the fronthaul ($D^{fr,t}$), midhaul ($D^{mid,t}$), and backhaul ($D^{b,t}$) links. Within each link, the transmission delay corresponds to the time needed to transmit all packets into the transmission medium, represented as $D=\frac{\alpha}{R}$, where $R$ indicates the data rate of the packet, and $\alpha$ denotes the mean packet size.
$D_{total}=D^{pro}+D^{tr},$ (14)
$D_{u,s}=\frac{L}{C}+\frac{\alpha}{R_{u,s}}.$ (15)
Therefore, the delay of UE $u$ in slice $s$, as derived from Equation 15, should be constrained to be below a specific threshold relevant to its respective slice, otulined below:
$D_{u,s}\leq D^{min}_{u,s}.$ (16)
The objective function is designed to optimize the overall network utility, which involves maximizing the weighted sum of the data rates achieved by UEs while ensuring compliance with QoS requirements. In this context, $w_{s}$ signifies the priority associated with individual service types and UEs. The intelligent RU is responsible for effectively managing the requirements of both network slices. Hence, we can mathematically express the problem in the following manner:
@@
$R_{u,s}\geq R_{s}^{max}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (21)
@@
$D_{u,s}\leq D^{min}_{u,s}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S},$ (23)

### v2

@@
Physical Resource Block
@@
Salar Nouri    Mojdeh Karbalaee Motalleb     Vahid Shah-Mansouri       Seyed Pooya Shariatpanahi
@@ Abstract
This paper introduces an innovative approach to the resource allocation problem, aiming to coordinate multiple independent x-applications (xAPPs) for network slicing and resource allocation in the Open Radio Access Network (O-RAN). Our approach maximizes the weighted throughput among user equipment (UE) and allocates physical resource blocks (PRBs). We prioritize two service types: enhanced Mobile Broadband and Ultra-Reliable Low-Latency Communication. Two xAPPs have been designed to achieve this: a power control xAPP for each UE and a PRB allocation xAPP. The method consists of a two-part training phase. The first part uses supervised learning with a Variational Autoencoder trained to regress the power transmission, UE association, and PRB allocation decisions, and the second part uses unsupervised learning with a contrastive loss approach to improve the generalization and robustness of the model. We evaluate the performance by comparing its results to those obtained from an exhaustive search and deep Q-network algorithms and reporting performance metrics for the regression task. The results demonstrate the superior efficiency of this approach in different scenarios among the service types, reaffirming its status as a more efficient and effective solution for network slicing problems compared to state-of-the-art methods. This innovative approach not only sets our research apart but also paves the way for exciting future advancements in resource allocation in O-RAN.
@@ I INTRODUCTION
Wireless communication has evolved from a mere convenience to an absolute necessity, driven by an insatiable demand for faster and more dependable connectivity between various quality of service (QoS) requirements [1]. The emergence of fifth generation (5G) technology represents a significant leap, offering the potential to support a wide range of applications, including augmented reality and autonomous vehicles [2]. However, the pursuit of wireless excellence does not conclude with 5G; it marks a stepping stone towards the next frontier— sixth generation (6G), the sixth generation of wireless technology. Poised on the horizon, 6G promises to redefine the boundaries of connectivity again.
@@
Emerging concepts within 6G, such as extreme Multiple-Input Multiple-Output and deep Receiver, promise to propel us further into uncharted technological realms, offering unprecedented data rates and levels of connectivity [3]. These advancements can potentially revolutionize our modes of communication and reshape the fabric of our interconnected world. Furthermore, the application of various machine learning (ML) techniques to automate resource allocation within the medium access control (MAC) layer and scheduler represents a burgeoning area of research within the 6G landscape [4]. These techniques aim to inject unparalleled intelligence into network management and optimization, laying the foundation for networks capable of dynamically adapting to user demands and environmental conditions. In this constantly evolving landscape, the roles of Open Radio Access Network (O-RAN) [5] and network slicing [6], which have reached remarkable milestones in 5G, continue to be pivotal. As we explore the possibilities offered by 6G and its transformative capabilities, it becomes increasingly clear that the evolution of wireless communication is an ongoing journey, establishing a robust foundation for the next era of wireless excellence.
@@
Resource allocation involves carefully distributing limited network resources among numerous users, applications, or services [7]. In 5G, the main goal is to support a wide range of services, including Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communications (URLLC), and massive Machine Type Communications (mMTC) [8]. Each service requires specific resource sets, including bandwidth, data rate, latency, power consumption, complexity, and reliability. Resource allocation within the 5G landscape needs to be flexible to meet these diverse demands while efficiently using network resources.
@@
Network slicing is an innovative concept that customizes virtual networks (slices) within a unified physical infrastructure to address specific service types and use cases [9]. In 5G, it efficiently allocates resources for eMBB and URLLC [10] services. This significance extends into the anticipated 6G era, where network slicing remains instrumental, facilitating resource allocation for diverse applications, including holographic communication and tactile internet [11, 12]. Its inherent scalability ensures adaptability to evolving workloads and user expectations, solidifying its role as a critical technology in the transition from 5G to 6G.
@@
This concept aligns with the architecture of O-RAN. O-RAN is based on open interfaces, open hardware, and open software and aims to provide a more flexible and cost-effective solution for wireless communication [7]. Implementing network slicing in O-RAN optimizes resource allocation and promotes multi-vendor interoperability and innovation.
@@
The core of O-RAN’s architecture consists of two fundamental components: Radio Units, responsible for physical layer functions, and Distributed Units, which manage baseband processing and control. These components work together, with the Central Unit (CU) serving as a centralized coordinator for network resources, facilitating dynamic resource allocation [13, 5]. The CU further comprises two distinct elements: the control plane (CP) and the user plane (UP). Adding to O-RAN’s adaptability and innovation are xApps (cross-applications), capable of spanning multiple network domains to optimize resource allocation for various services. These network segments include Fronthaul, Backhaul, and Midhaul, facilitating data transmission between components. Overseeing resource management and orchestration tasks is the Management and Orchestration (MANO) layer, ensuring the network’s efficiency. Network Functions Virtualization (NFV) decouples software from hardware, enhancing scalability. Furthermore, the Radio Access Network Intelligent Controller (RIC) plays a pivotal role in intelligent resource control, enabling dynamic allocation to meet diverse service requirements [5]. These components collectively define O-RAN’s transformative architecture, promoting openness, adaptability, and efficient resource utilization in modern wireless networks.
It is worth noting that O-RAN comprises two types of RIC: near-real-time (near-RT) RIC, which manages Radio Access Network (RAN) control and optimization, and non-real-time (non-RT) RIC. Within the near-RT RIC, xAPPs facilitate RAN control and optimization, while inside the non-RT RIC, rAPPs support it by providing guidance and employing ML models. Furthermore, both xAPPs and rAPPs can be implemented by third-party entities, allowing for a diverse ecosystem of applications within the O-RAN architecture [14].
@@
Using xAPPs leverages O-RAN network slicing to use resources efficiently and facilitate collaboration across services. However, this also introduces complexities in network management, as each xAPP needs to adjust the different QoS requirements of multiple services, posing a coordination challenge. Furthermore, cellular operators face the difficulty of establishing consistent connectivity for devices with varying QoS demands [15, 16]. Therefore, this study addresses the resource allocation problem within O-RAN using network slicing by considering the xAPPs. We propose a novel resource allocation method to effectively coordinate the numerous independent xAPPs to address network slicing and resource allocation challenges.
The use of ML algorithms in wireless networks has increased due to the growing complexity of these networks and the challenges of creating custom optimization models. Traditional approaches, such as convex/non-convex optimization, are struggling with the increasing complexity stemming from changes in network architecture, evolving performance requirements, and the growing number of devices. Current methods for network resource allocation that rely on learning-based techniques, like deep learning (DL) and deep reinforcement learning (DRL), often overlook the ability to learn resource distributions. Typically, these methods involve designing an agent for a specific task and training it from scratch, which reduces exploration efficiency and extends convergence times. Moreover, they require many training samples, leading to prolonged exploration phases.
The limited adaptability of traditional ML methods to new scenarios motivates us to explore an alternative approach capable of better generalization, more efficient data utilization, and resource distribution learning. To address these objectives, we propose a novel methodology that uses a self-supervised learning technique, specifically contrastive learning [17], in combination with a variational autoencoder (VAE) model, which is a type of generative model [18]. These models excel in handling uncertainty by learning distributions, offering a distribution for predictions instead of a singular deterministic output.
Moreover, their skill in handling incomplete or missing data allows them to fill in missing values based on learned patterns. VAEs also excel in semi-supervised learning setups, effectively using labeled and unlabeled data to improve predictive accuracy. Their generative abilities enable the creation of new data instances that resemble the training set, which is helpful in scenarios requiring synthetic data generation or increased resilience against noise and outliers.
@@
Furthermore, contrastive learning involves comparing different examples to acquire informative data representations. On the other hand, a VAE consists of an encoder and decoder that work together to learn the underlying probability distribution of data. By integrating contrastive learning with VAE, we can address the challenge of needing many training samples in ML. This approach can improve performance metrics, leading to faster convergence and better results while reducing the time and effort required for algorithm design and training.
In our proposed solution, x-Apps help to implement advanced resource allocation algorithms within the O-RAN framework. The algorithm is deployed as an x-App on the near-real-time RIC, enabling efficient and dynamic resource allocation adjustment based on real-time network conditions and user demands.
@@ I-A Contribution
This paper introduces a system model in the O-RAN architecture that provides support for two primary services, namely eMBB and URLLC. It focuses on resource allocation, with a particular emphasis on network slicing. The primary contributions of our paper can be summarized as follows:
  • Our research aims to address the resource allocation problem in the O-RAN architecture, explicitly focusing on optimizing weighted throughput by allocating resources such as transmission power and Physical Resource Blocks. We consider the unique QoS requirements and service priorities of two key 5G services: eMBB and URLLC. Compared to existing joint resource allocation schemes, our xAPP-based architecture enhances network management flexibility.
  • Our approach to optimizing resource allocation involves two main steps. Firstly, we simplify the problem by framing it as an optimization problem. Then, we introduce a novel machine learning algorithm that combines supervised learning with a VAE and unsupervised learning with contrastive loss to optimize resource allocation in O-RAN. It significantly improves the model’s generalization ability, making it a more practical solution for real-world applications.
  • While explicitly guaranteeing a global optimal solution is difficult, our method aims to achieve near-optimal solutions through evaluation with multiple benchmarks and significantly lower computational costs, making it a robust and practical solution for modern O-RAN systems. The method balances computational efficiency and scalability, making it suitable for real-time resource allocation in dynamic O-RAN environments. It is not limited to O-RAN and can be applied to future multi-vendor O-RAN architectures, leveraging semi-supervised generative model principles developed by our team.
@@ I-B Structure of Paper
The paper is structured as follows: Section II reviews related works. In Section III, we introduce the system model for the resource allocation problem in the O-RAN architecture and formally define the resource allocation problem as an optimization problem. Section IV provides a detailed description of the proposed resource allocation algorithm. Section V presents numerical results used to evaluate the algorithm’s performance. Section VI compares the computational cost of employed algorithms. Finally, Section VII offers our conclusion.
@@ II Related Works
As mentioned in Section I, cellular operators face significant challenges in ensuring reliable connections for devices with diverse QoS requirements. This challenge becomes more prominent in 5G and 6G networks due to their increased complexity and stringent QoS demands [16]. Recently, a growing focus has been on addressing the network slicing problem in cellular networks [16, 19, 20, 21, 22].
Feng et al. [16] proposed a method that employs network slicing and scheduling to allocate resources based on QoS requirements. Another study presented a resource allocation approach based on the M/M/1 queuing system for RAN slicing [19]. Nevertheless, practical scenarios often involve bursty traffic, overly optimistic, rendering the M/M/1 assumption of a Poisson distribution of packet arrivals at each gNodeB, which serves the eMBB and URLLC services. Consequently, the gNodeB cannot be considered an M/M/1 system, and its queue cannot be accurately modeled as a continuous-time Markov process. Additionally, in a study by Kazemifard et al., [20], an O-RAN system was investigated, assuming that the core network and RAN are co-located in the same data center. The authors conducted a theoretical analysis and employed a mathematical model to address minimum delay problems efficiently.
Various approaches have been proposed in the literature to address resource allocation challenges in wireless networks, including optimization techniques, game theory, and ML algorithms [23]. ML algorithms, in particular, have found application in power control and resource allocation within wireless networks.
@@
Similarly, Samarakoon et al. [30] introduced a federated learning algorithm to optimize power transmission and radio resource allocation in vehicular networks. Additionally, Elsayed et al. [31] proposed a reinforcement learning-based method for joint power and radio resource allocation in 5G networks.
@@ III System Model and Problem Formulation
In our study, we address the resource allocation problem using network slicing within O-RAN architecture, focusing on two primary slices: the URLLC slice and the eMBB slice. The URLLC slices ensure ultra-low latency and high reliability for critical applications such as autonomous driving and remote surgery. In contrast, eMBB slices focus on high data rates for mobile broadband. Both slices have different QoS requirements for network resource allocation [10].
Slicing the RAN into separate components for URLLC and eMBB optimizes the network for different traffic types and services, improving the efficient and flexible utilization of network resources. Key resources in O-RAN allocation include transmission power and PRBs. Traffic demands, often quantified by data rate and bandwidth, reflect the volume of data requiring transmission through the network. QoS requirements define the desired network performance based on factors such as data rate. The system model for our resource allocation problem with network slicing is illustrated in Figure 1. Table I describes all variables and parameters used in this study.
@@
In our study, we explore a deployment based on O-RAN, wherein each Base Station (BS), known as RU in O-RAN, supports two different slices: eMBB and URLLC. These slices are designed to meet the QoS requirements of User Equipments. Our analysis focuses on two key network functions, power control, and radio resource allocation, which are jointly considered. We represent the set of all RUs as $\mathcal{B}$ and the set of UEs as $\mathcal{U}$. Each BS $b$ serves two sets of UEs: $\mathcal{U}_{e,b}$ for eMBB and $\mathcal{U}_{u,b}$ for URLLC, leading to a total of $U$ UEs, where $U=U_{e,b}+U_{u,b}$. Each UE is exclusively associated with a single RU. The temporal and spectral layout is divided into $\mathcal{M}$ segments known as PRBs, which are the smallest unit of resources for allocation [32]. Each PRB covers a specific time slot in the temporal domain and a frequency range of f Hz in the spectral domain. We denote $S_{e}$ for eMBB slices and $S_{u}$ for URLLC slices. The total number of slices ($S$) will be $S=S_{e}+S_{u}$.
The UE association is indicated by a binary variable $\alpha_{u,b,s}$, which takes a value of 1 if UE $u$ is connected to RU $b$ in slice $s$, and 0 if not. Since each UE should be connected to only one RU, the following constraint must be met:
@@
We define the binary variable $\beta_{u,b,m,s}$ as the indicator for PRB allocation, taking the value of 1 when UE $u$ in slice $s$ is connected to PRB $m$ of RU $b$, and 0 otherwise.
@@
We denote the transmission power of PRB $m$ in RU $b$ in slice $s$ as $p_{u,b,m,s}$ for UE $u$, and the channel gain between RU $b$ and UE $u$ in slice $s$ as $h_{u,b,m,s}$. The rate achieved for a UE $u$ in slice $s$ by allocating PRB $m$ of RU $b$ is denoted by $R_{u,b,m,s}$, where $\eta_{u,b,m,s}$ represents the Signal to Interference & Noise Ratio (SINR). The calculation for $R_{u,b,m,s}$ is determined by:
@@
The total power transmitted by RU $b$ and the data rate for UE in the fronthaul link are represented as $P_{b}$ and $C_{b}$, respectively. Quantization noise ($\sigma_{1}^{2}$) can potentially impact the signal’s fidelity during transmission over the fronthaul, mainly when the original signal displays significant variations or spans a wide range of values. This noise may lead to inaccuracies in the transmitted signal, thereby influencing the QoS perceived by the UE. These values are determined as follows:
@@
The channel gain matrices in our model are generated using a Rayleigh fading process to simulate variable channel conditions typical in non-line-of-sight environments. The channel vector from the PRB $m$ in RU $b$ to the $u$ UE in the $s$ slice, denoted by $\mathbf{h}_{u,b,m,s}$, is modeled as $\mathbf{h}_{u,b,m,s}=\sqrt{\beta_{u,b,m,s}}\mathbf{g}_{u,b,m,s}$, where $\beta_{u,b,m,s}$ represents the large-scale fading coefficient, accounting for path loss and shadowing, and $\mathbf{g}_{u,b,m,s}\sim\mathcal{N}(0,N_{0}I_{D_{s}})$ is the fast and flat fading effects modeled as a circularly symmetric complex Gaussian random variable. To reflect the impact of interference and uncertainties, the channel gain matrix $\mathbf{H}$ is augmented with Gaussian noise, expressed as $\mathbf{H}_{noise}=\mathbf{H}+\mathbf{\epsilon}$, where $\mathbf{\epsilon}$ is a Gaussian noise matrix with zero mean and adjustable variance $\sigma^{2}$.
The total delay experienced by each UE is the sum of propagation and transmission delays. The overall propagation delay ($D^{pro}$) consists of the delays in the fronthaul ($D^{fr,p}$), midhaul ($D^{mid,p}$), and backhaul ($D^{b,p}$) links. In each of these links, the propagation delay is determined as the time it takes for a signal to traverse the distance, calculated as $D=\frac{L}{C}$, where $L$ is the link length and $C$ is the propagation speed. Similarly, the total transmission delay ($D^{tr}$) is the sum of transmission delays in the fronthaul ($D^{fr,t}$), midhaul ($D^{mid,t}$), and backhaul ($D^{b,t}$) links. Within each link, the transmission delay corresponds to the time needed to transmit all packets into the transmission medium, determined by $D=\frac{\mu}{R_{u,s}}$, where $R_{u,s}$ is the data rate for UE $u$ in slice $s$, and $\mu$ is the mean packet size. The total delay for UE $u$ in slice $s$ is $D_{u,s}=\frac{L}{C}+\frac{\mu}{R_{u,s}}$, and should be below a specific threshold ($D^{max}_{u,s}$), expressed as $D_{u,s}\leq D^{max}_{u,s}$.
The objective function is designed to optimize the overall network utility, which involves maximizing the weighted sum of the data rates achieved by UEs while ensuring satisfying the QoS requirements. In this context, $w_{s}$ indicates the priority associated with individual service types. The intelligent RU is responsible for effectively managing the requirements of both network slices. Hence, we can mathematically express the problem in the following manner:
@@
$R_{u,s}\geq R_{s}^{min}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (14)
@@
$D_{u,s}\leq D^{max}_{u,s}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S},$ (16)